OpenStudy (anonymous):

HEEEEELP MEH PLZ The non-profit organization you volunteer for is throwing a fundraiser cookout. You are in charge of buying the hamburgers, which cost $3.00 per pound and hotdogs, which cost $2.00 per pound. The meat budget you are given totals $600. The inequality 3x + 2y less than or greater to 600 represents the possible combinations of pounds of hamburgers (x) and hotdogs (y) you can buy. Which of the following represents a solution to the inequality?

OpenStudy (anonymous):

GRAPH

OpenStudy (anonymous):

WHICH OF THE FOLLOWING? WHAT FOLLOWING?

OpenStudy (anonymous):

Oops forgot to put the choices 200 pounds of hamburgers and 140 pounds of hotdogs 150 pounds of hamburgers and 60 pounds of hotdogs 100 pounds of hamburgers and 240 pounds of hotdogs 240 pounds of hamburgers and 40 pounds of hotdogs @.Gjallarhorn.

OpenStudy (anonymous):

@radar @mathstudent55 @Michele_Laino @misssunshinexxoxo

OpenStudy (michele_laino):

for example, if we use data from the first option, we get: \[3 \times 200 + 2 \times 140 = ...?\]

OpenStudy (anonymous):

880? @Michele_Laino

OpenStudy (michele_laino):

which is greater than 600, so the first option is not the right option

OpenStudy (michele_laino):

now, using data from your second option, we get: \[3 \times 150 + 2 \times 60 = ...?\]

OpenStudy (anonymous):

570? @Michele_Laino

OpenStudy (michele_laino):

Correct! Being 570 less than 600, what can you conclude?

OpenStudy (anonymous):

-30? im right?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

please wait a moment, I'm pondering....

OpenStudy (michele_laino):

I think that your problem asks for all pairs (x,y) such that the subsequent condition 3x+2y < 600 is checked

OpenStudy (michele_laino):

being 570 less than 600, then we can conclude that the pair (x=150, y=60) checks our condition above

OpenStudy (michele_laino):

for example if we use data from the third option we get: 3*100+ 2*240= 780 which is greater than600, so the third pair, doesn't check our condition

OpenStudy (michele_laino):

now, if we use data from the fourth option, we get: 3*240 + 2*40 = 720+80= 800> 600 so, also the pair (x=240, y=40) doesn't check our condition

OpenStudy (michele_laino):

so, what is the pair which checks our condition?

OpenStudy (anonymous):

:c

OpenStudy (anonymous):

@inkyvoyd Help?

OpenStudy (anonymous):

@Vocaloid >.<

OpenStudy (anonymous):

@brynndelyn

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@calcmathlete

OpenStudy (anonymous):

You're trying to satisfy \(3x+2y \le600\), where x is the pounds of hamburgers and y is the pounds of hot dogs. Basically, go through each of the answer choices and check to see which satisfy the equation. Does that make sense?

OpenStudy (anonymous):

*the inequality

OpenStudy (anonymous):

Ah I see. Thanks! @Calcmathlete

OpenStudy (anonymous):

No problem. :)